OpenStudy (shauna77777):

If v(t)=-t^2+4t+8 and s(2)=20 find the displacement s(t) and any point in time?

OpenStudy (shauna77777):

Do I find the antiderivative of v(t) and then plug in s(2) to find the constant? I'm not sure.

jimthompson5910 (jim_thompson5910):

hint: \[\Large s(t) = \int v(t)dt\]

jimthompson5910 (jim_thompson5910):

what is the antiderivative of v(t)?

OpenStudy (shauna77777):

I got s(t)=-(x^3/3)+(2x^2)+(8x)+C .. Would I plug into s(2) into this equation?

jimthompson5910 (jim_thompson5910):

`I got s(t)=-(x^3/3)+(2x^2)+(8x)+C ` replace every x with t

jimthompson5910 (jim_thompson5910):

\[\Large s(t) = -\frac{1}{3}t^3+2t^2+8t+C\]

jimthompson5910 (jim_thompson5910):

plug in t = 2 and that will have s(2) on the left side use the fact that s(2) = 20 to replace the "s(2)" with "20" then solve for C

OpenStudy (shauna77777):

Then s(t)=(-t^3/3)+2t^2+8t-1.33 would be the answer?

jimthompson5910 (jim_thompson5910):

C = -1.3333333333333... = -4/3

jimthompson5910 (jim_thompson5910):

So, \[\Large s(t) = -\frac{1}{3}t^3+2t^2+8t-\frac{4}{3}\]

OpenStudy (shauna77777):

Ah yes I should have put it in fraction form - thank you very much I was a little confused.

jimthompson5910 (jim_thompson5910):

I'm glad I could help out